# Clotho: Measuring Task-Specific Pre-Generation Test Adequacy for LLM Inputs

### v2

@@ Abstract
Software increasingly relies on the emergent capabilities of Large Language Models (LLMs), from natural language understanding to program analysis and generation. Yet testing them on specific tasks remains difficult and costly: many prompts lack ground truth, forcing reliance on human judgment, while existing uncertainty and adequacy measures typically require full inference. A key challenge is to assess input adequacy in a way that reflects the demands of the task, ideally before even generating any output. We introduce Clotho  [^1] Clotho, the Greek goddess of fate, spins the thread of human life. The name suggests that our technique operates at the start of LLM inference, the pre-generation stage., a task-specific, pre-generation adequacy measure that estimates input difficulty directly from hidden LLM states. Given a large pool of unlabelled inputs for a specific task, Clotho uses a Gaussian Mixture Model (GMM) to adaptively sample the most informative cases for human labelling. Based on this reference set the GMM can then rank unseen inputs by their likelihood of failure. In our empirical evaluation across eight benchmark tasks and three open-weight LLMs, Clotho can predict failures with a ROC-AUC of 0.716, after labelling reference sets that are on average only 5.4% of inputs. It does so without generating any outputs, thereby reducing costs compared to existing uncertainty measures. Comparison of Clotho and post-generation uncertainty measures shows that the two approaches complement each other. Crucially, we show that adequacy scores learnt from open-weight LLMs transfer effectively to proprietary models, extending the applicability of the approach. When prioritising test inputs for proprietary models, Clotho increases the average number of failing inputs from 18.7 to 42.5 out of 100, compared to random prioritisation.
1 Introduction
@@
The emergent capabilities of Large Language Models (LLMs) have enabled new types of software systems, from domain-specific natural language understanding Kim et al. (2023) to autonomous coding and testing agents Zhang et al. (2024); Yang et al. (2024); Yoon et al. (2024); Rondon et al. (2025). As LLMs move beyond text generation into specific inference and reasoning tasks and see more widespread adoption, testing them becomes more critical and more costly. Yet the very qualities that make LLMs powerful also make them difficult to validate: their inner workings are opaque, and for many tasks we rely on them to solve, no ground truth is immediately available. Our goal is to improve task-specific LLM testing by prioritising which inputs to run and label first, ideally exposing failures earlier and with less effort.
@@
Prior work on test adequacy in Deep Neural Networks (DNNs) have focused on structural coverage Pei et al. (2017); Ma et al. (2018) and distributional deviation from training data Kim et al. (2019); Kim and Yoo (2021). These test adequacy metrics enable us to prioritise manual labelling for new inputs Kim et al. (2020), or to guide automated input generation Tian et al. (2018). However, these approaches do not translate well to LLMs. First, LLM pre-training corpora are vast and heterogeneous Gao et al. (2020); Laurençon et al. (2023), making it infeasible to assess coverage or distributional density. Second, what matters in practice is not pre-training behaviour, it is rather how the model performs on a specific task as described by a fixed prompt [^2] Technically a prompt is a prompt template describing a task, into which specific inputs are then inserted..
@@
Since adequacy cannot be judged against pre-training data, we can turn to signals derived from the model itself. Transformer-based LLMs Vaswani et al. (2017) encode input as embeddings, process it through stacked attention and feed-forward layers, and decode responses token by token. Within this structure, several post-hoc uncertainty measures have been proposed: semantic entropy to detect hallucinations Farquhar et al. (2024), self-consistency as a proxy for reliability Wang et al. (2023), and measures like token-level confidence and inference variability to flag input risk Huang et al. (2025). While these metrics are informative, they are all post-generation. They exploit signals available only after output is produced, but thus require running full inference, often multiple times due to output non-determinism, which increases both computation time and cost. This limits their practicality for large-scale testing.
@@
To tackle these challenges, we introduce Clotho, a task-scoped adequacy measure designed around two key elements. First, it embeds testing in an active learning loop Ren et al. (2021) that incrementally expands a small seed set of labelled inputs, reducing reliance on costly ground truth by focusing human effort where it matters most. Second, it operates pre-generation, using only input embeddings of an LLM to estimate adequacy before any output is produced. This avoids the overhead of full inference while still identifying inputs likely to reveal failures.
@@
Concretely, Clotho constructs a Gaussian Mixture Model over the Last-token Input Hidden States (LIHS) of passing inputs in the labelled seed set. This surrogate model (i.e., predicting uncertainty in place of the target LLM) ranks unseen inputs by their likelihood-based “surprise”, in the spirit of Surprise Adequacy Kim et al. (2019, 2022), with low-likelihood inputs treated as more challenging. The reference set is expanded iteratively: new labelling targets are chosen where the existing distribution is ambiguous, and also where they add diversity, so that human effort focuses on the most informative cases. To align with reference set expansion, the GMM adapts its number of components and dimensionality guided by cross-validation and estimates of active component usage. While pre-generation signals require hidden states, we show that scoring models trained on small open-weight LLMs transfer effectively to closed-weight models (e.g., GPT, Claude, Gemini), enabling cost-effective testing even with API-only access.
@@
2 Exploratory Study: Transformer Hidden States under Task-specific Prompting
Clotho’s core assumption is that both the type of task and the relative difficulty of each input for a task are reflected in the internal hidden states within the Transformer architecture, even before any output token is produced. To verify the validity of these assumptions, several key questions need to be answered: (1) How do the Transformer internal states represent different tasks when the model is prompted with specific instructions? (2) Can we reliably distinguish potentially failing inputs from straightforward passing inputs based solely on Transformer internal states? (3) Which layers of the Transformer architecture provide the most useful hint for characterising input difficulty? In this section, to explore the feasibility of leveraging hidden states to derive test adequacy metrics, we present preliminary studies that aim to answer these questions.
@@ 2 Exploratory Study: Transformer Hidden States under Task-specific Prompting
Figure 1: Visualisations on Llama 8B internal state space with inputs for all studied tasks
@@
Figure 2: Pass/Fail Separation of SPELL-CHK Inputs in Projected PCA Space Fitted on Passing Inputs
@@
2.1 Modelling Representative Cases in Task-specific Internal State Subspace
Figures 1(a) and 1(b) plot activation vectors from a hidden layer of Llama3.1 8B across inputs from eight natural and programming language tasks studied in this paper (see Section 4.2). We reduced the vectors to two dimensions with t-SNE Maaten and Hinton (2008). In Figure 1(a), colours mark tasks; in Figure 1(b), green and red mark passing and failing inputs. It shows that the eight tasks form well-separated clusters in the input hidden state space. More importantly, even in the reduced dimension, we observe that passing and failing inputs tend to separate well, suggesting that internal hidden states contain information that can distinguish easier inputs from the more challenging ones.
@@ 2.1 Modelling Representative Cases in Task-specific Internal State Subspace
We further examined how failures align with the distribution of passing inputs: we fitted Principal Component Analysis (PCA) Abdi and Williams (2010) on the activations of passing inputs from the SPELL-CHK task, and then transformed all inputs from the same SPELL-CHK task, including failing ones, into the fitted space (Figure 1(c)). We observe a clear pattern where failing inputs deviate from the dense region of passing inputs. However, we note that this distributional deviation is task-specific. Figure 1(d) presents the same SPELL-CHK inputs, but mapped into the PCA space fitted with the passing inputs of a different task, TOPIC-CLS. In this cross-task projection, the separation between passing and failing inputs become notably weaker compared to the within-task projection. From this observation, we posit that distributional deviation in hidden state space of LLMs can indeed guide test adequacy, but only when analysed in a task-specific way. PCA was used here because t-SNE does not transfer across different data samples.
@@
Finally, Figure 1(a) also shows that some tasks yield multi-modal distributions, with several distinct clusters instead of a single dense region. This means that out-of-distribution detection cannot be simply reduced to measuring distance from the centre of one unimodal distribution. It calls instead for models that can capture multi-modality Kim and Yoo (2021).
@@
Figure 3: Input hidden states of Llama 8B from two prompted tasks (one per line) for varying depths of layers.
2.2 Extracting Internal Hidden Vectors from Transformer Architecture
@@
Layer choice is less obvious. Prior work shows that middle-to-late layers capture semantic abstractions more effectively, for example, in hallucination detection Kossen et al. (2024); Azaria and Mitchell (2023) and adversarial input identification Zhou et al. (2025). We verify this in our setting by examining hidden states of the final input token across all layers in Llama3.1 8B (32 layers in total), using one natural language task (SPELL-CHK) and one programming task (SYN-BUG). Figure 3 illustrates the results for a subset of the layers, visualised with t-SNE, after first reducing the dimensionality to 10 using PCA for computational efficiency. Early layers show little distinction between passing and failing cases. In middle and later layers, however, separation emerges, with inputs clustering more distinctly by outcome. The effect seems strongest in the latter half of the model’s depth.
@@
3 Approach
Clotho formulates the input-level test adequacy as the degree of distributional deviation Kim et al. (2019, 2022). To align the reference distribution with the given task, without relying on LLM training data, Clotho adaptively models the distribution of inputs the target LLM can successfully handle. This is achieved by fitting a Gaussian Mixture Model (GMM) to hidden state embeddings (LIHS) of passing inputs, expanded through strategic sampling. This section describes details of the approach.
3.1 Measuring Test Adequacy as Distributional Deviation
Given a labelled reference set of inputs, Clotho estimates probability density of passing inputs in the space of LLM hidden states. More concretely, we extract LIHS of passing inputs and apply Gaussian Mixture Model to capture their distribution. The use of GMM is motivated by the fact that the distribution of LIHS may not be unimodal (see Section 2.1).
@@ 3.1 Measuring Test Adequacy as Distributional Deviation
3 while $|\mathcal{R}|<N$ do
4   Adapt $\mathcal{M}_{t}$ with best representation dim. and # of mixture components w.r.t. $\mathcal{R}$ ;
5 $\mathcal{R}_{\text{pass}}\leftarrow\{x\in\mathcal{R}\mid\text{label}(x)=\text{pass}\}$ ;
6   Fit GMM $\mathcal{M}_{t}$ on $\mathcal{R}_{\text{pass}}$ ;
7
8 for $x\in\mathcal{U}$ do
9    $s_{\text{exploit}}(x)\leftarrow$ entropy of component responsibilities under $\mathcal{M}_{t}$ ;
10
11
12 $B_{\text{explore}}\leftarrow\lfloor\alpha B\rfloor$, $B_{\text{exploit}}\leftarrow B-B_{\text{explore}}$ ;
13
14 $C_{\text{exploit}}\leftarrow$ the $B_{\text{exploit}}$ inputs with the highest $s_{\text{exploit}}$ values ;
15 $C_{\text{explore}}\leftarrow\emptyset$ ;
16  t for $j=1$ to $B_{\text{explore}}$ do
17      For each $x\in\mathcal{U}\setminus(C_{\text{exploit}}\cup C_{\text{explore}})$,
18      compute $d(x)=\min_{y\in\mathcal{R}\cup C_{\text{exploit}}\cup C_{\text{explore}}}\text{Dist}(x,y)$ ;
19      Pick $x^{*}=\arg\max_{x}d(x)$ ;
20    $C_{\text{explore}}\leftarrow C_{\text{explore}}\cup\{x^{*}\}$ ;
21
22 $C\leftarrow C_{\text{exploit}}\cup C_{\text{explore}}$ ;
23
24  Execute and obtain human labels for inputs in $C$ ;
25 $\mathcal{R}\leftarrow\mathcal{R}\cup C$, $\mathcal{U}\leftarrow\mathcal{U}\setminus C$, $t\leftarrow t+1$ ;
26
@@ 24  Execute and obtain human labels for inputs in $C$ ;
In the main loop (lines 7-19), Clotho expands the reference set $\mathcal{R}$. We cast this as an active learning process Cohn et al. (1996), where inputs are iteratively selected for labelling. Each round targets inputs that maximise information gain using two complementary strategies: exploitation, which sharpens the boundaries of the known hidden state space, and exploration, which extends coverage into under-represented regions. To balance these goals, Clotho divides the batch into $B_{exploit}$ and $B_{explore}$ (line 9). Exploitation is guided by the GMM’s component assignment entropy on unlabelled inputs (lines 8 and 10). High entropy signals uncertainty about which component an input belongs to, so labelling the most uncertain inputs helps refine the model’s understanding of the distribution. Exploration, in contrast, is driven by diversity: Clotho selects inputs with the greatest minimum Euclidean distance from those already in $\mathcal{R}$ (line 14), following the greedy max-min strategy of Adaptive Random Testing (ART) Chen et al. (2005). Larger distances ensure broader coverage of the hidden state space. Since no prior knowledge of the latent space is available, Clotho weights exploitation and exploration equally, setting $\alpha=0.5$: we may consider adaptive fine-tuning in the future.
Using Algorithm 1, Clotho produces a Gaussian Mixture Model (GMM), $\mathcal{M}_{T}$, that estimates the probability density of hidden states from passing inputs for a given task:
$\mathcal{L}_{\text{pass}}(x)=p_{\theta}\!\left(h(x)\right),\quad h(x)\text{: hidden state representation (LIH)}$
The density $p_{\theta}\!\left(h(x)\right)$ captures how plausible an input is relative to the reference distribution of LIHS from passing cases: higher values indicate that $x$ falls within regions commonly associated with correct outputs. From this density, we can compute Likelihood-based Surprise Adequacy (LSA) Kim et al. (2019), defined as the negative log likelihood:
@@
3.2 Adaptive Modelling of Passing Input Distribution via GMM
In Algorithm 1, the model $\mathcal{M}_{t}$ is iteratively updated as the reference set expands. To improve model accuracy, we adapt two key parameters during this process: the dimensionality of the LIH representations used by the GMM and the number of mixture components in the model. Both adaptations are performed at line 4 of Algorithm 1.
3.2.1 Dimensionality Selection

### v3

@@ Abstract.
Software increasingly relies on the emergent capabilities of Large Language Models (LLMs), from natural language understanding to program analysis and generation. Yet testing them on specific tasks remains difficult and costly: many prompts lack ground truths, forcing reliance on human judgments, while existing test adequacy measures typically rely on output uncertainty and thus are only available after full inference. A key challenge is to assess how useful a test input is in a way that reflects the demands of the task, ideally before even generating any output. We introduce Clotho  [^1] Clotho, the Greek goddess of fate, spins the thread of human life. The name suggests that our technique operates at the start of LLM inference, the pre-generation stage., a task-specific, pre-generation test adequacy measure that estimates input difficulty directly from LLM hidden states. Given a large pool of unlabelled inputs for a specific task, Clotho uses a Gaussian Mixture Model (GMM) to adaptively sample the most informative cases for human labelling. Based on this reference set the GMM can then rank unseen inputs by their likelihood of failure. In our empirical evaluation across eight benchmark tasks and three open-weight LLMs, Clotho can predict failures with a ROC-AUC of 0.716, after labelling reference sets that are on average only 5.4% of inputs. It does so without generating any outputs, thereby significantly reducing LLM execution costs compared to output-based uncertainty or confidence measures. Comparison of Clotho and these post-generation adequacy measures shows that the two approaches complement each other. Crucially, we show that adequacy scores learnt from open-weight LLMs transfer effectively to proprietary models, extending the applicability of the approach. When prioritising test inputs for proprietary models, Clotho increases the average number of failing inputs from 18.7 to 42.5 out of 100, compared to random prioritisation.
Software Testing, Software Verification, Test Adequacy, Test Prioritisation, LLM-based Application
copyright: ccdoi: 10.1145/3797114journalyear: 2026journal: PACMSEjournalvolume: 3journalnumber: FSEarticle: FSE086publicationmonth: 7submissionid: fse26maina-p3213-pccs: Software and its engineering Software testing and debuggingccs: Software and its engineering Empirical software validationccs: Software and its engineering Operational analysis
1. Introduction
@@
The emergent capabilities of Large Language Models (LLMs) have enabled new types of software systems, from domain-specific natural language understanding (Kim et al., 2023) to autonomous coding and testing agents (Zhang et al., 2024a; Yang et al., 2024; Yoon et al., 2024; Rondon et al., 2025). As LLMs move beyond text generation into specific inference and reasoning tasks and see more widespread adoption, testing them becomes more critical and more costly. Yet the very qualities that make LLMs powerful also make them difficult to validate: their inner workings are opaque, and for many tasks we rely on them to solve, no ground truth is immediately available. Our goal is to improve task-specific LLM testing by prioritising which inputs to run and label first, ideally exposing failures earlier and with less effort.
@@
Prior work on test adequacy in Deep Neural Networks (DNNs) have focused on structural coverage (Pei et al., 2017; Ma et al., 2018) and distributional deviation from training data (Kim et al., 2019; Kim and Yoo, 2021). These test adequacy metrics enable us to prioritise manual labelling for new inputs (Kim et al., 2020), or to guide automated input generation (Tian et al., 2018). However, these approaches do not translate well to LLMs. First, LLM pre-training corpora are vast and heterogeneous (Gao et al., 2020; Laurençon et al., 2023), making it infeasible to assess coverage or distributional density. Second, what matters in practice is not pre-training behaviour, it is rather how the model performs on a specific task as described by a fixed prompt [^2] Technically a prompt is a prompt template describing a task, into which specific inputs are then inserted..
@@
Since adequacy cannot be judged against pre-training data, we can turn to signals derived from the model itself. Transformer-based LLMs (Vaswani et al., 2017) encode input as embeddings, process it through stacked attention and feed-forward layers, and decode responses token by token. Within this structure, several post-hoc uncertainty measures have been proposed: semantic entropy to detect hallucinations (Farquhar et al., 2024), self-consistency as a proxy for reliability (Wang et al., 2023), and measures like token-level confidence and inference variability to flag input risk (Huang et al., 2025). While these metrics are informative, they are all post-generation. They exploit signals available only after output is produced, but thus require running full inference, often multiple times due to output non-determinism, which increases both computation time and cost. This limits their practicality for large-scale testing.
@@
To tackle these challenges, we introduce Clotho, a task-scoped adequacy measure designed around two key elements. First, it embeds testing in an active learning loop (Ren et al., 2021) that incrementally expands a small seed set of labelled inputs, reducing reliance on costly ground truth by focusing human effort where it matters most. Second, it operates pre-generation, using only input embeddings of an LLM to estimate adequacy before any output is produced. This avoids the overhead of full inference while still identifying inputs likely to reveal failures.
@@
Concretely, Clotho constructs a Gaussian Mixture Model (GMM) over the hidden states from the last token of passing inputs in labelled seed set. This surrogate model (i.e., predicting uncertainty in place of the target LLM) ranks unseen inputs by their likelihood-based “surprise”, in the spirit of Surprise Adequacy (Kim et al., 2019, 2022), with low-likelihood inputs treated as more challenging. The reference set is expanded iteratively: new labelling targets are chosen where the existing distribution is ambiguous, and also where they add diversity, so that human effort focuses on the most informative cases. To align with reference set expansion, the GMM adapts its number of components and dimensionality guided by cross-validation and estimates of active component usage. While pre-generation signals require hidden states, we show that scoring models trained on small open-weight LLMs transfer effectively to closed-weight models (e.g., GPT, Claude, Gemini), enabling cost-effective testing even with API-only access.
@@
2. Exploratory Study: Transformer Hidden States under Task-specific Prompting
Clotho’s core assumption is that both the type of task and the relative difficulty of each input for a task are reflected in the internal hidden states within the Transformer architecture, even before any output token is produced. To verify the validity of these assumptions, several key questions need to be answered: (1) How do the Transformer internal states represent different tasks when the model is prompted with specific instructions? (2) Can we reliably distinguish potentially failing inputs from straightforward passing inputs based solely on Transformer internal states? (3) Which layers of the Transformer architecture provide the most useful hint for characterising input difficulty?
In this section, to explore the feasibility of leveraging hidden states to derive test adequacy metrics, we present preliminary studies that aim to answer these questions.
@@ 2. Exploratory Study: Transformer Hidden States under Task-specific Prompting
Figure 1. Visualisation on Llama 8B internal state space using inputs for all studied tasks
@@
Figure 2. Pass/fail separation of SPELL-CHK inputs in projected PCA space of passing inputs
@@
2.1. Modelling Representative Cases in Task-specific Internal State Subspace
Figures 1(a) and 1(b) plot activation vectors from a hidden layer of Llama 3.1 8B for inputs from eight natural and programming language tasks studied in this paper (see Section 4.2). We reduced the vectors to two dimensions using t-SNE (Maaten and Hinton, 2008). In Figure 1(a), colours mark tasks; in Figure 1(b), colours represent the proportion of correct outputs across ten repeated generations, with greener shades indicating more correct outputs and redder shades indicating more incorrect outputs. It shows that the eight tasks form well-separated clusters in the internal representation space of inputs. To ensure that this structure is not an artefact of t-SNE exaggeration, we compute the silhouette score in the 10-dimensional PCA space, obtaining a value of 0.9579, which indicates exceptionally strong clustering. Consistently, intra-cluster cosine distances are very small (mean = 0.0265, max = 0.0490), whereas inter-cluster distances are substantially larger (mean = 1.1272). [^3] The smallest inter-cluster distance (0.2617) occurs between GH-TYPO and SPELL-CHK, which is intuitively reasonable given their shared focus on misspellings and typos.
More importantly, Figure 1(b) shows a noticeable separation between green and red regions even in the highly reduced 2D space, suggesting that internal input representations may contain information that distinguishes easier inputs from the more challenging ones. We further examined how input distributions vary depending on their pass/fail outcomes: we fitted Principal Component Analysis (PCA) (Abdi and Williams, 2010) on the activations of inputs for which the majority of repeated generations are correct (hereafter referred to as passing inputs) from the SPELL-CHK task, and then transformed all inputs from the same SPELL-CHK task, including failing ones, into the fitted space (Figure 1(c)). We observe a clear pattern where failing inputs deviate from the dense region of passing inputs. However, we note that this distributional deviation is task-specific.
@@ 2.1. Modelling Representative Cases in Task-specific Internal State Subspace
Figure 1(d) presents the same SPELL-CHK inputs, but mapped into the PCA space fitted with the passing inputs of a different task, TOPIC-CLS. In this cross-task projection, the separation between passing and failing inputs become notably weaker compared to the within-task projection. Examining this effect through the Fisher Discriminant Ratio (FDR) (Fisher, 1936) between passing and failing groups, we find that within-task PCA consistently yields stronger discriminability than cross-task PCA, including SPELL-CHK (0.736 vs. 0.314), ODD-ADD (0.593 vs. 0.246), and JSON-FIX (0.101 vs. 0.015). From this observation, we posit that distributional deviation in hidden state space of LLMs can indeed provide a useful signal for assessing test adequacy, but only when analysed in a task-specific way. PCA was used here because t-SNE does not transfer across different data samples.
@@
Finally, Figure 1(a) also shows that some tasks yield multi-modal distributions, with several distinct clusters instead of a single dense region. This means that out-of-distribution detection cannot be simply reduced to measuring distance from the centre of one unimodal distribution. It calls instead for models that can capture multi-modality (Kim and Yoo, 2021).
@@
Figure 3. Input hidden states of Llama 8B from two prompted tasks (one per line) for varying depths of layers.
2.2. Extracting Internal Hidden Vectors from Transformer Architecture
@@
Layer choice is less obvious. Prior work shows that middle-to-late layers capture semantic abstractions more effectively, for example, in hallucination detection (Kossen et al., 2024; Azaria and Mitchell, 2023) and adversarial input identification (Zhou et al., 2025). We verify this in our setting by examining hidden states of the final input token across all layers in Llama3.1 8B (32 layers in total), using one natural language task (SPELL-CHK) and one programming task (SYN-BUG). Figure 3 illustrates the results for a subset of the layers, visualised with t-SNE, after first reducing the dimensionality to 10 using PCA for computational efficiency. Early layers show little distinction between passing and failing cases. In middle and later layers, however, separation emerges, with inputs clustering more distinctly by outcome. The effect seems strongest in the latter half of the model’s depth.
@@
3. Approach
Clotho formulates the input-level test adequacy as the degree of distributional deviation (Kim et al., 2019, 2022). To align the reference distribution with the given task, without relying on LLM training data, Clotho adaptively models the distribution of inputs the target LLM can successfully handle. This is achieved by fitting a Gaussian Mixture Model (GMM) to hidden state embeddings (LIHS) of passing inputs, expanded through strategic sampling. This section describes details of the approach.
3.1. Measuring Test Adequacy as Distributional Deviation
Given a labelled reference set of inputs, Clotho estimates probability density of passing inputs in the space of LLM hidden states. More concretely, we extract LIHS of passing inputs and apply GMM to capture their distribution. The use of GMM is motivated by the fact that the distribution of LIHS may not be unimodal (see Section 2.1).
@@ 3.1. Measuring Test Adequacy as Distributional Deviation
while $|\mathcal{R}|<N$ do
   Adapt $\mathcal{M}_{t}$ with best representation dim. and # of mixture components w.r.t. $\mathcal{R}$ ;
 $\mathcal{R}_{\text{pass}}\leftarrow\{x\in\mathcal{R}\mid\text{label}(x)=\text{pass}\}$ ;
   Fit GMM $\mathcal{M}_{t}$ on $\mathcal{R}_{\text{pass}}$ ;
 for $x\in\mathcal{U}$ do
    $s_{\text{exploit}}(x)\leftarrow$ entropy of component responsibilities under $\mathcal{M}_{t}$ ;
   end for
 $B_{\text{explore}}\leftarrow\lfloor\alpha B\rfloor$, $B_{\text{exploit}}\leftarrow B-B_{\text{explore}}$ ;
 $C_{\text{exploit}}\leftarrow$ the $B_{\text{exploit}}$ inputs with the highest $s_{\text{exploit}}$ values ;
 $C_{\text{explore}}\leftarrow\emptyset$ ;
 for $j=1$ to $B_{\text{explore}}$ do
      For each $x\in\mathcal{U}\setminus(C_{\text{exploit}}\cup C_{\text{explore}})$,
      compute $d(x)=\min_{y\in\mathcal{R}\cup C_{\text{exploit}}\cup C_{\text{explore}}}\text{Dist}(x,y)$ ;
      Pick $x^{*}=\arg\max_{x}d(x)$ ;
    $C_{\text{explore}}\leftarrow C_{\text{explore}}\cup\{x^{*}\}$ ;
   end for
 $C\leftarrow C_{\text{exploit}}\cup C_{\text{explore}}$ ;
  Execute and obtain human labels for inputs in $C$ ;
 $\mathcal{R}\leftarrow\mathcal{R}\cup C$, $\mathcal{U}\leftarrow\mathcal{U}\setminus C$, $t\leftarrow t+1$ ;
end while
@@
In the main loop (lines 7-19), Clotho expands the reference set $\mathcal{R}$. We cast this as an active learning process (Cohn et al., 1996), where inputs are iteratively selected for labelling. Each round targets inputs that maximise information gain using two complementary strategies: exploitation, which sharpens the boundaries of the known hidden state space, and exploration, which extends coverage into under-represented regions. To balance these goals, Clotho divides the batch into $B_{exploit}$ and $B_{explore}$ (line 9). Exploitation is guided by the GMM’s component assignment entropy on unlabelled inputs (lines 8 and 10). High entropy signals uncertainty about which component an input belongs to, so labelling the most uncertain inputs helps refine the model’s understanding of the distribution. Exploration, in contrast, is driven by diversity: Clotho selects inputs with the greatest minimum Euclidean distance from those already in $\mathcal{R}$ (line 14), following the greedy max-min strategy of Adaptive Random Testing (ART) (Chen et al., 2005). Larger distances ensure broader coverage of the hidden state space. Since no prior knowledge of the latent space is available, Clotho weights exploitation and exploration equally, setting $\alpha=0.5$: we may consider adaptive fine-tuning in the future.
Using Algorithm 1, Clotho produces a Gaussian Mixture Model, $\mathcal{M}_{T}$, that estimates the probability density of hidden states from passing inputs for a given task:
$\mathcal{L}_{\text{pass}}(x)=p_{\theta}\!\left(h(x)\right),\quad h(x)\text{: hidden state representation (LIHS)}$
The density $p_{\theta}\!\left(h(x)\right)$ captures how plausible an input is relative to the reference distribution of LIHS from passing cases: higher values indicate that $x$ falls within regions commonly associated with correct outputs. From this density, we can compute Likelihood-based Surprise Adequacy (LSA) (Kim et al., 2019), defined as the negative log likelihood:
@@
3.2. Adaptive Modelling of Passing Input Distribution via GMM
In Algorithm 1, the model $\mathcal{M}_{t}$ is iteratively updated as the reference set expands. To improve model accuracy, we adapt two key parameters during this process: the dimensionality of the LIHS representations used by the GMM and the number of mixture components in the model. Both adaptations are performed at line 4 of Algorithm 1.
3.2.1. Dimensionality Selection